# ThinkEdit: Interpretable Weight Editing to Mitigate Overly Short Thinking in Reasoning Models

### v2

@@ 4.2 Identify Short Reasoning Attention Heads
$\overline{C}^{h}\;=\;\frac{1}{|\mathcal{D}_{\text{short}}|}\sum_{i\in\mathcal{D}_{\text{short}}}\left(\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}C^{h}(i,t)\right).$ (4)
@@
$\overline{C}^{h}_{\textrm{short}}\;=\;\left\langle\overline{C}^{h},-\hat{v}^{\text{attn}}_{\ell}\right\rangle.$ (5)
@@ 4.3 Editing Short Reasoning Heads
We introduce how ThinkEdit effectively removes the short reasoning direction from the output projection matrices of the "short reasoning heads". Specifically, we identify the top 2% of attention heads with the largest $\overline{C}^{h}_{\text{short}}$ values (as defined in Section 4.2), marking them as short reasoning heads. Let $W_{o}^{h_{\ell}}\in\mathbb{R}^{d_{h}\times d}$ be the output projection matrix of head $h$ in layer $\ell$, and let $-\hat{v}^{\text{attn}}_{\ell}\in\mathbb{R}^{d}$ denote the short reasoning direction at layer $\ell$. We then update $W_{o}^{h_{\ell}}$ via:
$W_{o}^{h_{\ell}}\;\leftarrow\;W_{o}^{h_{\ell}}\Bigl{(}I\;-\;\bigl{(}-\hat{v}^{\text{attn}}_{\ell}\bigr{)}\bigl{(}-\hat{v}^{\text{attn}}_{\ell}\bigr{)}^{\!\top}\Bigr{)},$ (6)
@@ 4.4 Performance of Reasoning Models after ThinkEdit
  • GSM8K [2]: A test set of 1,319 grade-school-level math word problems.
  • MMLU Elementary Math [6]: A subset of 378 elementary school math questions from the MMLU benchmark.
  • MATH-Level1: A collection of 437 easy (Level 1) problems drawn from the MATH dataset [7].
  • MATH-Level5: The most challenging subset of the MATH dataset, containing 1,324 difficult problems.
  • MATH-500 [13]: A curated set of 500 high-quality, diverse math problems designed to assess advanced mathematical reasoning.
@@
Table 1 reports the overall accuracy (in %) before and after applying ThinkEdit. Across all math benchmarks, we observe consistent improvements in accuracy. Notably, the deepseek-distill-qwen-1.5B model shows a substantial gain on the MMLU Elementary Math subset. Manual inspection reveals that the unedited model occasionally ignores the multiple-choice format, outputting a direct answer instead of selecting from the given options. In contrast, the edited model adheres to the instructions more reliably. This suggests that removing short reasoning components from specific attention heads may not only enhance reasoning quality but also improve instruction-following behavior. On the more challenging MATH-Level5 and MATH-500 datasets, the accuracy gains are more modest but still positive, suggesting that while editing short-reasoning heads has a stronger impact on simpler problems, it might still provide meaningful improvements even for harder tasks that require longer and more complex reasoning chains.
Model GSM8K MMLU Elem. Math MATH-Level1 MATH-Level5 MATH-500
ThinkEdit $\mathbf{93.50\pm 0.31}$ $\mathbf{96.53\pm 0.54}$ $\mathbf{96.50\pm 0.46}$ $\mathbf{91.15\pm 0.59}$ $\mathbf{91.78\pm 0.58}$
ThinkEdit $\mathbf{88.97\pm 0.78}$ $\mathbf{96.08\pm 0.86}$ $\mathbf{94.12\pm 0.47}$ $\mathbf{85.91\pm 0.48}$ $\mathbf{87.60\pm 0.81}$
ThinkEdit $\mathbf{83.34\pm 0.79}$ $\mathbf{86.24\pm 1.12}$ $\mathbf{93.89\pm 0.76}$ $74.94\pm 0.85$ $\mathbf{82.74\pm 0.77}$
@@
Model GSM8K MMLU Elem. Math MATH-Level1 MATH-Level5 MATH-500
ThinkEdit 96.62 / 96.03 / 96.12 96.11 / 96.22 / 96.27 100.00 / 99.77 / 98.85 95.76 / 97.65 / 98.07 89.60 / 92.60 / 94.70
ThinkEdit 97.08 / 95.27 / 93.95 97.78 / 98.65 / 97.87 100.00 / 99.30 / 98.62 95.61 / 96.89 / 97.12 92.80 / 93.60 / 94.40
ThinkEdit 92.46 / 92.37 / 92.05 77.22 / 80.54 / 79.73 96.19 / 95.81 / 97.36 93.79 / 95.83 / 95.80 92.80 / 94.40 / 94.90
@@
Finally, Table 3 reports the average reasoning length among the top 5%, 10%, and 20% shortest responses. We observe that ThinkEdit consistently increases the reasoning length in these short-answer scenarios, suggesting that the intervention discourages excessively short reasoning, and therefore leads to higher accuracy as shown in Table 2. Interestingly, despite the local increase, the average reasoning length across all responses changes by +3.5% for deepseek-qwen-14b, +3.6% for deepseek-llama3-8b, and -6.7% for deepseek-qwen-1.5b. As a result, the overall net change in reasoning length across all models is a slight decrease of 0.5%. This suggests that ThinkEdit selectively extends short reasoning traces without unnecessarily prolonging already long ones.
Model GSM8K MMLU Elem. Math MATH-Level1 MATH-Level5 MATH-500
ThinkEdit 95.4 / 106.3 / 120.2 79.1 / 87.1 / 98.7 125.1 / 150.2 / 243.4 698.5 / 906.6 / 1157.2 270.2 / 492.6 / 733.3
ThinkEdit 93.2 / 106.9 / 127.4 396.5 / 464.2 / 543.2 137.4 / 173.3 / 277.1 791.2 / 954.8 / 1185.1 305.2 / 506.3 / 737.6
ThinkEdit 97.2 / 109.4 / 126.3 75.9 / 85.0 / 99.5 127.9 / 174.1 / 416.4 818.0 / 984.5 / 1214.3 435.0 / 612.9 / 800.6
In summary, we identify a targeted subset of attention heads that drive overly short reasoning and introduce ThinkEdit as an effective intervention to address this issue. Appendix A.3 provides several examples demonstrating how ThinkEdit enhances reasoning quality.
@@ 5 Related Works
Recent advancements in reasoning models have enhanced LLMs’ problem-solving capabilities across mathematics, coding, and science. OpenAI’s o1 [9] marks a paradigm shift toward deliberate reasoning by leveraging Reinforcement Learning (RL) to refine its problem-solving strategies. With a "chain-of-thought" mechanism, o1 generates structured reasoning before responding, enabling superior performance on complex tasks like AIME. This integration of RL enhances its adaptability, establishing o1 as a significant milestone in AI-driven reasoning. DeepSeek-r1 [5] provides a cost-efficient alternative to o1, demonstrating that pure RL can effectively enhance reasoning capabilities. It introduces Group Relative Policy Optimization (GRPO) [18], an improvement that eliminates the need for a reward model, enabling more efficient RL training. Additionally, DeepSeek has distilled its R1 model into multiple sizes of Llama3 and Qwen models, broadening access to advanced reasoning models.
@@
Controllable text generation has been extensively explored across various domains [12], with methods generally classified into training-time and inference-time control. These approaches aim to steer LLMs toward generating text with specific attributes while preserving fluency and coherence. Training-time control is primarily achieved through fine-tuning [27, 24] or reinforcement learning [16, 17], leveraging diverse datasets to shape model behavior. Inference-time control encompasses prompt engineering [19, 11], representation engineering [20, 30, 10], interpretable neuron intervention [22, 23], and decoding-time interventions [3], allowing flexible and efficient control without model retraining.
In this work, we focus on the representation engineering paradigm, which involves modifying the internal representations of language models, to investigate how reasoning length is embedded within model representations. Specifically, we introduce a linear "reasoning-length direction" in the representation space to examine its impact on reasoning capabilities in LLMs.
Attention heads and MLP-layer neurons intervention
A growing body of research explores the role of attention heads and neurons within the Multi-Layer Perceptron (MLP) layers in shaping language model behavior. Studies such as [29, 28, 1] examine how safety mechanisms are embedded in well-aligned models to defend against harmful prompts and jailbreak attacks [31, 14, 21]. Findings indicate that a small subset of attention heads and MLP neurons play a critical role in safety alignments. Fine-tuning these specific components has been shown to significantly enhance model robustness against adversarial prompts. Similarly, research on hallucination mitigation has investigated the contributions of attention heads and MLP neurons. [8] demonstrates that filtering out unreliable attention heads can reduce erroneous generations, while [26] finds that activating subject-knowledge neurons helps maintain factual consistency.
These insights suggest that targeted interventions on key components of LLMs can be leveraged to steer model behavior, enhance factuality, and mitigate vulnerabilities. In our work, we investigate how these components relate to reasoning processes, examining their impact on structured problem-solving and logical inference.
@@ 6 Conclusion
In this work, we first identified overly short reasoning as a common failure mode in Deepseek-distilled reasoning models, often resulting in reduced accuracy. To understand how reasoning length is controlled, we analyzed the model’s hidden representations and uncovered a latent direction linked to reasoning length, with middle layers playing a key role. Building on this, we pinpointed a small set of attention heads that drive short reasoning, and propose ThinkEdit to mitigate the issue, leading to significant accuracy gains for short reasoning outputs (+5.44%), along with an overall improvement across multiple math benchmarks (+2.43%).
Future work could extend these techniques to achieve a more fine-grained understanding of how LLMs control reasoning length or explore other attributes, providing deeper insights into the mechanistic interpretability of reasoning models. Additionally, steering reasoning length could improve inference budget control, balancing reasoning depth and cost. This offers a promising direction for future research on efficient and adaptable models.
@@ Appendix A Appendix
A.1 Layerwise Analysis of Steering along Reasoning Length Direction

### v3

@@ 4.2 Identify Short Reasoning Attention Heads
$\overline{C}^{h}=\frac{1}{|\mathcal{D}_{\text{short}}|}\sum_{i\in\mathcal{D}_{\text{short}}}\left(\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}C^{h}(i,t)\right).\vspace{-5pt}$ (4)
@@
$\overline{C}^{h}_{\textrm{short}}\;=\;\left\langle\overline{C}^{h},-\hat{v}^{\text{attn}}_{\ell}\right\rangle.\vspace{-5pt}$ (5)
@@ 4.3 Editing Short Reasoning Heads
We introduce how ThinkEdit effectively removes the short reasoning direction from the output projection matrices of the "short reasoning heads". Specifically, we identify the top 4% of attention heads with the largest $\overline{C}^{h}_{\text{short}}$ values (as defined in Section 4.2), marking them as short reasoning heads. Let $W_{o}^{h_{\ell}}\in\mathbb{R}^{d_{h}\times d}$ be the output projection matrix of head $h$ in layer $\ell$, and let $-\hat{v}^{\text{attn}}_{\ell}\in\mathbb{R}^{d}$ denote the short reasoning direction at layer $\ell$. We then update $W_{o}^{h_{\ell}}$ via:
$W_{o}^{h_{\ell}}\;\leftarrow\;W_{o}^{h_{\ell}}\Bigl{(}I\;-\;\bigl{(}-\hat{v}^{\text{attn}}_{\ell}\bigr{)}\bigl{(}-\hat{v}^{\text{attn}}_{\ell}\bigr{)}^{\!\top}\Bigr{)},\vspace{-5pt}$ (6)
@@
Model GSM8K MMLU Elem. Math MATH-Level1 MATH-Level5 MATH-500
ThinkEdit (4%) $\mathbf{93.78\pm 0.50}$ $\mathbf{96.56\pm 0.84}$ $\mathbf{96.36\pm 0.52}$ $\mathbf{91.03\pm 0.44}$ $\mathbf{91.92\pm 0.63}$
ThinkEdit (4%) $\mathbf{89.44\pm 0.55}$ $\mathbf{96.19\pm 0.73}$ $\mathbf{94.44\pm 0.31}$ $\mathbf{86.49\pm 0.54}$ $\mathbf{88.06\pm 1.09}$
ThinkEdit (4%) $\mathbf{84.56\pm 0.79}$ $\mathbf{90.66\pm 0.97}$ $\mathbf{93.66\pm 0.62}$ $75.05\pm 0.82$ $\mathbf{82.24\pm 0.89}$
Model GSM8K MMLU Elem. Math MATH-Level1 MATH-Level5 MATH-500
ThinkEdit (4%) 96.31 / 96.18 / 96.77 97.78 / 95.14 / 96.53 99.52 / 99.53 / 98.62 96.67 / 97.88 / 98.11 91.20 / 93.20 / 95.00
ThinkEdit (4%) 96.31 / 95.50 / 94.68 97.78 / 97.57 / 97.60 99.05 / 99.07 / 97.82 95.76 / 97.42 / 97.46 95.60 / 93.80 / 95.40
ThinkEdit (4%) 92.62 / 92.90 / 92.32 87.78 / 88.11 / 88.67 95.71 / 95.58 / 96.44 95.15 / 96.59 / 97.27 90.80 / 92.00 / 94.20
@@ 4.4 Performance of Reasoning Models after ThinkEdit
  • GSM8K Cobbe et al. (2021): A test set of 1,319 grade-school-level math word problems.
  • MMLU Elementary Math Hendrycks et al. (2021a): A subset of 378 elementary school math questions from the MMLU benchmark.
  • MATH-Level1: A collection of 437 easy (Level 1) problems drawn from the MATH dataset Hendrycks et al. (2021b).
  • MATH-Level5: The most challenging subset of the MATH dataset with 1,324 problems.
  • MATH-500 Lightman et al. (2023): A curated set of 500 high-quality math problems designed to assess advanced mathematical reasoning.
@@
Table 1 reports the overall accuracy (in %) before and after applying ThinkEdit. Across all math benchmarks, we observe consistent improvements in accuracy. Notably, the deepseek-distill-qwen-1.5B model shows a substantial gain on the MMLU Elementary Math subset. Manual inspection reveals that the unedited model occasionally ignores the multiple-choice format, leading to wrong answers. In contrast, the edited model adheres to the instructions more reliably. This suggests that ThinkEdit may not only enhance reasoning quality but also improve instruction-following behavior. On the more challenging MATH-Level5 and MATH-500 datasets, the accuracy gains are more modest but still positive, suggesting that while editing short-reasoning heads has a stronger impact on simpler problems, it might still provide meaningful improvements even for harder tasks that require longer and more complex reasoning chains.
@@
We analyze how ThinkEdit affects reasoning length in Appendix A.3. It modestly increases the length of the shortest responses (Table 3), helping to address overly brief reasoning. However, as shown in Table 4, the overall reasoning length remains largely stable across datasets, with a net change of -0.27% across all models and datasets.
We also explore different editing percentages and compare our approach to simply appending “Wait” to prompt longer reasoning; detailed results are provided in Appendix A.4. Additionally, results for deepseek-distill-qwen-32B are reported in Appendix A.5. Finally, We provide several concrete examples illustrating how ThinkEdit enhances reasoning quality in Appendix A.7.
@@ 5 Related Works
Recent advances in reasoning models have significantly improved the problem-solving abilities of LLMs in domains such as mathematics, coding, and science. OpenAI’s o1 Jaech et al. (2024) represents a major shift toward deliberate reasoning by employing reinforcement learning (RL) to refine its strategies. By generating explicit "Thinking" steps before producing answers, o1 achieves strong performance on complex tasks. As a more cost-efficient alternative, DeepSeek-r1 Guo et al. (2025) demonstrates that pure RL can also effectively enhance reasoning. It introduces Group Relative Policy Optimization (GRPO) Shao et al. (2024), a novel method that eliminates the need for a separate reward model, enabling more efficient RL training.
@@
Controllable text generation has been explored across various domains Liang et al. (2024), with methods generally classified into training-time and inference-time control. These approaches aim to steer LLMs toward generating text with specific attributes while preserving fluency and coherence. Training-time control is achieved through fine-tuning Zeldes et al. (2020); Wei et al. (2022) or reinforcement learning Ouyang et al. (2022); Rafailov et al. (2023), leveraging diverse datasets to shape model behavior. Inference-time control encompasses prompt engineering Shin et al. (2020); Li and Liang (2021), representation engineering Subramani et al. (2022); Zou et al. (2023a); Konen et al. (2024), interpretable neuron intervention Sun et al. (2024b, c), and decoding-time interventions Dathathri et al. (2020), allowing flexible and efficient control without retraining.
In this work, we focus on the representation engineering paradigm to investigate how reasoning length is embedded within model representations. Specifically, we introduce a linear "reasoning-length direction" in the representation space to examine its impact on reasoning capabilities.
Attention heads and MLP neurons intervention.
A growing body of research explores the role of attention heads and neurons within the Multi-Layer Perceptron (MLP) layers in shaping language model behavior. Studies such as Zhou et al. (2025); Zhao et al. (2025); Chen et al. (2024) examine how safety mechanisms are embedded in well-aligned models to defend against harmful prompts and jailbreak attacks Zou et al. (2023b); Liu et al. (2024); Sun et al. (2024a). Findings indicate that a small subset of attention heads and MLP neurons play a critical role in safety alignments. Similarly, research on hallucination mitigation has investigated the contributions of attention heads and MLP neurons. Ho et al. (2025) demonstrates that filtering out unreliable attention heads can reduce erroneous generations, while Yu et al. (2024) finds that activating subject-knowledge neurons helps maintain factual consistency.
In our work, we investigate how attention heads relate to reasoning processes, examining their impact on the reasoning length and quality.
@@ 6 Conclusion
In this work, we first identified overly short reasoning as a common failure mode in Deepseek-distilled reasoning models. To understand how reasoning length is controlled, we analyzed the model’s hidden representations and uncovered a latent direction linked to reasoning length. Building on this, we pinpointed 4% of attention heads that drive short reasoning, and propose ThinkEdit to mitigate the issue, leading to significant accuracy gains for short reasoning outputs (+6.39%), along with an overall improvement (+3.34%) across multiple math benchmarks.
Limitations
A limitation of our work is that ThinkEdit primarily improves model performance by addressing cases of overly short reasoning. For reasoning models that already tend to produce sufficiently long or verbose outputs, the benefits of ThinkEdit may be limited. Nonetheless, our study provides valuable insights by highlighting the often-overlooked issue of overly brief reasoning and examining how reasoning length is represented within the model’s hidden states. This opens an important research direction for advancing the interpretability of reasoning models by linking internal representations to observable reasoning behaviors.
@@ Appendix A Appendix
A.1 Gobal Steering with the MLP-based Direction $v_{\ell}^{\text{mlp}}$
Figure 5 replicates the global steering analysis using the MLP-based direction $v_{\ell}^{\text{mlp}}$. The observed trends closely mirror those from attention-based steering: increasing $\alpha$ extends reasoning length across both datasets, and the effect on accuracy is model- and dataset-dependent. On GSM8K, larger models benefit from longer reasoning, while smaller models degrade. On MATH-Level5, overly long reasoning may hurt performance, despite consistent control over CoT length. These results show that both attention and MLP directions capture similar reasoning-length attributes.
Figure 5: Global steering with the reasoning length direction extracted from MLPs. The trend is similar as steering with attention-based directions.
A.2 Layerwise Analysis of Steering along Reasoning Length Direction